# Cost Efficient Fairness Audit Under Partial Feedback

### v1

@@ 7 Conclusion and Future Work
In this work we studied the fairness audit problem under the partial feedback model. We introduced a novel cost model to capture real-world concerns of exploring additional data. With this cost model, we studied a non-parametric (blackbox) and a mixture of exponential family settings for data distributions. Our framework also combined historical partial feedback data along with online samples. We provided a complete characterization of the blackbox setting with matching upper and lower bounds upto logarithmic factors. In the mixture model setting, we leveraged techniques from learning with truncated samples and MAP oracles to provide an algorithm with significantly lower cost than the blackbox case. We complement our theoretical findings with empirical results that demonstrate the cost efficiency of our proposed algorithms. In this work, we established the connection between partial feedback model with learning from truncated samples which we believe to be a promising research direction. Another interesting future work would be to explore the audit cost complexity of unknown classifiers from a known hypothesis class, e.g., threshold classifiers. Moreover, one can also ask what are other data assumptions under which algorithms with lower audit cost can be designed.
@@ Appendix A Appendix for Blackbox Model
Recall that the quantities of interest are $\beta\coloneqq\mathbb{P}[f=0],\beta_{a}\coloneqq\mathbb{P}[f=0\mid A=a],p_{y,a}\coloneqq\mathbb{P}[f=1,Y=y,A=a]$, $q_{y,a}\coloneqq\mathbb{P}[Y=y,A=a]$, $\gamma_{ij}\coloneqq\mathbb{P}[f=i,Y=j]$ for $i,j\in\{0,1\}$, $p_{y\mid a}\coloneqq\mathbb{P}[f=1,Y=y\mid A=a]$ and $q_{y\mid a}\coloneqq\mathbb{P}[Y=y\mid A=a]$ for all $y\in\{0,1\}$ and $a\in{\mathcal{A}}$. Note that $p_{y,a}=p_{y\mid a}\cdot\mathbb{P}[A=a]$, therefore $p_{y,a}\leq p_{y\mid a}$, and similarly, $q_{y,a}\leq q_{y\mid a}$. We also have $p_{0,a}\leq p_{0}$ for all $a\in{\mathcal{A}}$.
@@ A.1.1 Proof of Theorem 1
$\texttt{cost}\leq\tilde{O}\left(\sum_{a\in\mathcal{A}}\sum_{y\in\{0,1\}}\frac{c_{feat}\beta+c_{lab}\gamma_{00}}{q_{y,a}\varepsilon^{2}}\right)~.$
Moreover, the number of true labels requested by this policy is at most $\tilde{O}(\sum_{a\in{\mathcal{A}}}\sum_{y\in\{0,1\}}\frac{\beta}{q_{y,a}\varepsilon^{2}})$.
@@
Finally, the expected number of individuals for whom $h=1$ is $N\cdot\mathbb{P}[h=1]=N\cdot\mathbb{P}[f=0]=N\beta$. This is the actual number of individuals for whom the algorithm requests true labels. Hence, the number of true labels requested by this algorithm is at most
$N\cdot\beta\leq\sum_{a\in\mathcal{A}}\sum_{y\in\{0,1\}}\frac{624\beta\log(8\lvert\mathcal{A}\rvert/\delta)}{\varepsilon^{2}q_{y,a}}~.$
Further, the audit cost of observing the true labels of $N$ individuals is $N\cdot(c_{feat}\mathbb{P}[f=0]+c_{lab}\mathbb{P}[f=0,Y=0])=N\cdot(c_{feat}\beta+c_{lab}\gamma_{00})$. Hence, the audit cost of the algorithm is bounded by
$\texttt{cost}\leq\sum_{a\in\mathcal{A}}\sum_{y\in\{0,1\}}\frac{624(c_{feat}\beta+c_{lab}\gamma_{00})\log(8\lvert\mathcal{A}\rvert/\delta)}{\varepsilon^{2}q_{y,a}}~.$
@@
Suppose the true hypothesis is $\mathsf{UNFAIR}$. Let $y^{*}\in\{0,1\}$ and $a_{1},a_{2}\in\mathcal{A},a_{1}\neq a_{2}$ be such that $\lvert\mathbb{P}[f=1\mid Y=y^{*},A=a_{1}]-\mathbb{P}[f=1\mid Y=y^{*},A=a_{2}]\rvert=\left\lvert\frac{p_{y^{*},a_{1}}}{q_{y^{*},a_{1}}}-\frac{p_{y^{*},a_{2}}}{q_{y^{*},a_{2}}}\right\rvert>\varepsilon$. Then, we have,
@@
$>\varepsilon-\frac{\varepsilon}{2}$ ($\left\lvert\frac{p_{y^{*},a_{1}}}{q_{y^{*},a_{1}}}-\frac{p_{y^{*},a_{2}}}{q_{y^{*},a_{2}}}\right\rvert>\varepsilon$)
$=\frac{\varepsilon}{2}~.$
@@
$\lvert\mathbb{P}[f=1\mid Y=y,A=a]-\mathbb{P}[f=1\mid Y=y,A=a^{\prime}]\rvert=\left\lvert\frac{p_{y,a}}{q_{y,a}}-\frac{p_{y,a^{\prime}}}{q_{y,a^{\prime}}}\right\rvert=0~.$
@@
$=0+\frac{\varepsilon}{2}$ ($\left\lvert\frac{p_{y,a}}{q_{y,a}}-\frac{p_{y,a^{\prime}}}{q_{y,a^{\prime}}}\right\rvert=0$)
$=\frac{\varepsilon}{2}~.$
In particular, we have $\widehat{\Delta}=\max_{y\in\{0,1\}}\max_{a,a^{\prime}\in{\mathcal{A}}}\left\lvert\frac{\widehat{p}_{y,a}}{\widehat{q}_{y,a}}-\frac{\widehat{p}_{y,a^{\prime}}}{\widehat{q}_{y,a^{\prime}}}\right\rvert\leq\frac{\varepsilon}{2}$. Hence, the algorithm outputs $\mathsf{FAIR}$ implying that the algorithm’s output is correct.
@@ A.2 Performance of RS-Audit in Blackbox Model
$\texttt{cost}\leq\tilde{O}\left(\sum_{a\in\mathcal{A}}\sum_{y\in\{0,1\}}\frac{c_{feat}\beta_{a}+c_{lab}(q_{0\mid a}-p_{0\mid a})}{q_{y\mid a}\varepsilon^{2}}\right)~.$
Moreover, the number of true labels requested by this policy is at most $\tilde{O}(\sum_{a\in{\mathcal{A}}}\sum_{y\in\{0,1\}}\frac{\beta_{a}}{q_{y\mid a}\varepsilon^{2}})$.
@@
Note that $N_{y\mid a}$ samples are collected by the algorithm under the conditional distribution $\mathbb{P}[Y\mid A=a]$. Therefore, the expected number of samples actually requested by the algorithm is $N_{y\mid a}\cdot\mathbb{P}[f=0\mid A=a]=N_{y\mid a}\beta_{a}$. Similarly, the audit cost of $N_{y\mid a}$ samples is $N_{y\mid a}(c_{feat}\mathbb{P}[f=0\mid A=a]+c_{lab}\mathbb{P}[f=0,Y=0\mid A=a])=N_{y\mid a}(c_{feat}\beta_{a}+c_{lab}(q_{0\mid a}-p_{0\mid a}))$. Therefore, the expected number of samples requested by the algorithm is at most
$\sum_{y\in\{0,1\}}\sum_{a\in{\mathcal{A}}}N_{y\mid a}\cdot\beta_{a}\leq\sum_{a\in\mathcal{A}}\sum_{y\in\{0,1\}}\frac{624\beta_{a}\log(8\lvert\mathcal{A}\rvert/\delta)}{\varepsilon^{2}q_{y\mid a}}~.$
@@
$\texttt{cost}=\sum_{y\in\{0,1\}}\sum_{a\in{\mathcal{A}}}N_{y\mid a}(c_{feat}\beta_{a}+c_{lab}(q_{0\mid a}-p_{0\mid a}))\leq\sum_{a\in\mathcal{A}}\sum_{y\in\{0,1\}}\frac{624(c_{feat}\beta_{a}+c_{lab}(q_{0\mid a}-p_{0\mid a}))\log(8\lvert\mathcal{A}\rvert/\delta)}{\varepsilon^{2}q_{y\mid a}}~.$
@@
Suppose the true hypothesis is $\mathsf{UNFAIR}$. Let $y^{*}\in\{0,1\}$ and $a_{1},a_{2}\in\mathcal{A},a_{1}\neq a_{2}$ be such that $\lvert\mathbb{P}[f=1\mid Y=y^{*},A=a_{1}]-\mathbb{P}[f=1\mid Y=y^{*},A=a_{2}]\rvert=\left\lvert\frac{p_{y^{*}\mid a_{1}}}{q_{y^{*}\mid a_{1}}}-\frac{p_{y^{*}\mid a_{2}}}{q_{y^{*}\mid a_{2}}}\right\rvert>\varepsilon$. Then, we have,
@@
$>\varepsilon-\frac{\varepsilon}{2}$ ($\left\lvert\frac{p_{y^{*}\mid a_{1}}}{q_{y^{*}\mid a_{1}}}-\frac{p_{y^{*}\mid a_{2}}}{q_{y^{*}\mid a_{2}}}\right\rvert>\varepsilon$)
$=\frac{\varepsilon}{2}~.$
@@
$\lvert\mathbb{P}[f=1\mid Y=y,A=a]-\mathbb{P}[f=1\mid Y=y,A=a^{\prime}]\rvert=\left\lvert\frac{p_{y\mid a}}{q_{y\mid a}}-\frac{p_{y\mid a^{\prime}}}{q_{y\mid a^{\prime}}}\right\rvert=0~.$
@@
$=0+\frac{\varepsilon}{2}$ ($\left\lvert\frac{p_{y\mid a}}{q_{y\mid a}}-\frac{p_{y\mid a^{\prime}}}{q_{y\mid a^{\prime}}}\right\rvert=0$)
$=\frac{\varepsilon}{2}~.$
@@ A.3 Lower Bound for the Blackbox Model
Under the blackbox setting, for any $\varepsilon\in(0,\nicefrac{{1}}{{4}})$ and $\delta\in(0,1)$, there exists a family of instances characterized by $p_{y\mid a}$, $q_{y\mid a}$, and $\beta_{a}$ for $y\in\{0,1\}$ and $a\in{\mathcal{A}}$, even with $\lvert\mathcal{A}\rvert=2$, such that any algorithm with finite expected running time that succeeds in the $(\varepsilon,\delta)$-fairness test must request at least $\widetilde{\Omega}\left(\sum_{a\in\mathcal{A}}\sum_{y\in\{0,1\}}\frac{\beta_{a}}{q_{y\mid a}\varepsilon^{2}}\right)$ true labels in expectation.
@@
Let $p,q\in(0,1)$ be constants to be decided later. The classifier $f$ is given as follows. Under the $\mathsf{FAIR}$ hypothesis, given $Y=1$ and $A=a$, $f\sim\mathrm{Ber}(\frac{1}{2})$ and given $Y=0$ and $A=a$, $f\sim\mathrm{Ber}(\frac{p}{2(1-q)})$ for both $a\in\{0,1\}$. On the other hand, in the $\mathsf{UNFAIR}$ hypothesis, the distribution of $f$ is as specified in Table 1.
@@
From Table 2, it can be seen that given $A=a$, $\mathbb{P}_{\mathsf{FAIR}}[f=1\mid A=a]=\mathbb{P}_{\mathsf{UNFAIR}}[f=1\mid A=a]$. Moreover, the $\mathbb{P}_{\mathsf{FAIR}}[Y=1\mid f=1,A=a]=\mathbb{P}_{\mathsf{UNFAIR}}[Y=1\mid f=1,A=a]$. Hence, $D_{KL}(\mathbb{P}_{\mathsf{FAIR}}[Y\mid f=1,A=a]\parallel\mathbb{P}_{\mathsf{UNFAIR}}[Y\mid f=1,A=a])=0$ Finally, $\mathbb{P}_{\mathsf{FAIR}}[Y=1\mid f=0,A=1]=\frac{q}{2(1-\frac{p+q}{2})}$ and $\mathbb{P}_{\mathsf{UNFAIR}}[Y=1\mid f=0,A=1]=\frac{q(1+8\varepsilon)}{2(1-\frac{p+q}{2})}$.
We set $\varepsilon\in(0,\frac{1}{4})$, $p\in(0,\frac{1}{2})$ and $q\in(0,\frac{1}{2})$ so that $q(1+4\varepsilon)\leq 1$. It is easy to verify that the equalized odds difference under the $\mathsf{FAIR}$ hypothesis is 0 (intuitively, the probabilities of the classifier is the same for both the groups $a\in\{0,1\}$). The equalized odds disparity gap of the $\mathsf{UNFAIR}$ hypothesis is more than $\varepsilon$ as can be seen below.
@@
$\geq\frac{1}{2}\exp\left(-64\cdot\mathbb{E}_{\mathsf{FAIR}}[L_{\tau}]\cdot\varepsilon^{2}\cdot\frac{q}{2(1-\frac{p+q}{2})}\right)$ (Lemma 12, $\frac{q}{2-p-q}\leq\frac{1}{2}$ and $\frac{8q\varepsilon}{2-p-q}\leq 1$ for $q\in(0,\frac{1}{2})$ and $\varepsilon\in(0,\frac{1}{4})$)
Therefore, $\mathbb{E}_{\mathsf{FAIR}}[L_{\tau}]>\frac{(1-\frac{p+q}{2})\log(1/4\delta)}{32q\varepsilon^{2}}$. Moreover, since Lemma 11 is symmetric, we also have $\mathbb{E}_{\mathsf{UNFAIR}}[L_{\tau}]>\frac{(1-\frac{p+q}{2})\log(1/4\delta)}{64q\varepsilon^{2}}$ using Lemma 12. Hence, $\min\{\mathbb{E}_{\mathsf{FAIR}}[L_{\tau}],\mathbb{E}_{\mathsf{UNFAIR}}[L_{\tau}]\}>\frac{(1-\frac{p+q}{2})\log(1/4\delta)}{64q\varepsilon^{2}}$. Therefore, under both the hypothesis, the algorithm must request $\Omega\left(\frac{\mathbb{P}[f=0\mid A=1]\log(1/\delta)}{\mathbb{P}[Y=1\mid A=1]\varepsilon^{2}}\right)$ samples in expectation to be able to output the correct hypothesis.

### v2

@@ 7 Conclusion and Future Work
In this work we initiated the study of fairness audit under the partial feedback model and introduced a novel cost model to capture real-world concerns of exploring additional data. With this cost model, we studied blackbox and mixture of exponential family models of data distributions. Our framework combines historical partial feedback data along with online samples. We provided a complete characterization of the blackbox setting with matching upper and lower bounds. In the mixture model setting, we leveraged techniques from learning with truncated samples and approximate MAP oracles to design an algorithm with significantly lower cost than the blackbox case. We complement our theoretical findings with empirical results that demonstrate the cost efficiency of our proposed algorithms. An interesting future work is to explore the audit cost complexity of unknown classifiers from a known hypothesis class, e.g., threshold classifiers. Moreover, one can also study non-uniform cost models.
Acknowledgments and Disclosure of Funding
Nirjhar Das acknowledges the support of the Walmart Center for Tech Excellence (CSR WMGT-23-0001).
@@ Appendix A Appendix for Blackbox Model
Recall that the quantities of interest are $p_{y,a}\coloneqq\mathbb{P}[f=1,Y=y,A=a]$, $q_{y,a}\coloneqq\mathbb{P}[Y=y,A=a]$, $p_{y\mid a}\coloneqq\mathbb{P}[f=1,Y=y\mid A=a]$, and $q_{y\mid a}\coloneqq\mathbb{P}[Y=y\mid A=a]$ for all $y\in\{0,1\}$ and $a\in{\mathcal{A}}$. Note that $p_{y,a}=p_{y\mid a}\cdot\mathbb{P}[A=a]$, therefore $p_{y,a}\leq p_{y\mid a}$, and similarly, $q_{y,a}\leq q_{y\mid a}$.
@@ A.1.1 Proof of Theorem 1
$\texttt{cost}\leq\widetilde{O}\left(\sum_{a\in\mathcal{A}}\sum_{y\in\{0,1\}}\frac{c_{feat}\mathbb{P}[f=0]+c_{lab}\mathbb{P}[f=0,Y=0]}{q_{y,a}\varepsilon^{2}}\right)~.$
Moreover, the number of true labels requested by this policy is at most $\tilde{O}(\sum_{a\in{\mathcal{A}}}\sum_{y\in\{0,1\}}\frac{\mathbb{P}[f=0]}{q_{y,a}\varepsilon^{2}})$.
@@
Finally, the expected number of individuals for whom $h=1$ is $N\cdot\mathbb{P}[h=1]=N\cdot\mathbb{P}[f=0]$. This is the actual number of individuals for whom the algorithm requests true labels. Hence, the number of true labels requested by this algorithm is at most
$N\cdot\mathbb{P}[f=0]\leq\sum_{a\in\mathcal{A}}\sum_{y\in\{0,1\}}\frac{624\mathbb{P}[f=0]\log(8\lvert\mathcal{A}\rvert/\delta)}{\varepsilon^{2}q_{y,a}}~.$
Further, the audit cost of observing the true labels of $N$ individuals is $N\cdot(c_{feat}\mathbb{P}[f=0]+c_{lab}\mathbb{P}[f=0,Y=0])$. Hence, the audit cost of the algorithm is bounded by
$\texttt{cost}\leq\sum_{a\in\mathcal{A}}\sum_{y\in\{0,1\}}\frac{624(c_{feat}\mathbb{P}[f=0]+c_{lab}\mathbb{P}[f=0,Y=0])\log(8\lvert\mathcal{A}\rvert/\delta)}{\varepsilon^{2}q_{y,a}}~.$
@@
Suppose the true hypothesis is $\mathsf{UNFAIR}$. Let $y^{*}\in\{0,1\}$ and $a_{1},a_{2}\in\mathcal{A},a_{1}\neq a_{2}$ be such that $\lvert\mathbb{P}[f=1\mid Y=y^{*},A=a_{1}]-\mathbb{P}[f=1\mid Y=y^{*},A=a_{2}]\rvert=\left\lvert\frac{p_{y^{*},a_{1}}}{q_{y^{*},a_{1}}}-\frac{p_{y^{*},a_{2}}}{q_{y^{*},a_{2}}}\right\rvert>\gamma+\varepsilon$. Then, we have,
@@
$>\gamma+\varepsilon-\frac{\varepsilon}{2}$ ($\left\lvert\frac{p_{y^{*},a_{1}}}{q_{y^{*},a_{1}}}-\frac{p_{y^{*},a_{2}}}{q_{y^{*},a_{2}}}\right\rvert>\gamma+\varepsilon$)
$=\gamma+\frac{\varepsilon}{2}~.$
@@
$\lvert\mathbb{P}[f=1\mid Y=y,A=a]-\mathbb{P}[f=1\mid Y=y,A=a^{\prime}]\rvert=\left\lvert\frac{p_{y,a}}{q_{y,a}}-\frac{p_{y,a^{\prime}}}{q_{y,a^{\prime}}}\right\rvert\leq\gamma~.$
@@
$\leq\gamma+\frac{\varepsilon}{2}~.$ ($\left\lvert\frac{p_{y,a}}{q_{y,a}}-\frac{p_{y,a^{\prime}}}{q_{y,a^{\prime}}}\right\rvert\leq\gamma$)
In particular, we have $\widehat{\Delta}=\max_{y\in\{0,1\}}\max_{a,a^{\prime}\in{\mathcal{A}}}\left\lvert\frac{\widehat{p}_{y,a}}{\widehat{q}_{y,a}}-\frac{\widehat{p}_{y,a^{\prime}}}{\widehat{q}_{y,a^{\prime}}}\right\rvert\leq\gamma+\frac{\varepsilon}{2}$. Hence, the algorithm outputs $\mathsf{FAIR}$ implying that the algorithm’s output is correct.
@@ A.2 Performance of RS-Audit in Blackbox Model
$\texttt{cost}\leq\widetilde{O}\left(\sum_{a\in\mathcal{A}}\sum_{y\in\{0,1\}}\frac{c_{feat}\mathbb{P}[f=0\mid A=a]+c_{lab}\mathbb{P}[f=0,Y=0\mid A=a]}{q_{y\mid a}\varepsilon^{2}}\right)~.$
Moreover, the number of true labels requested by this policy is at most $\tilde{O}(\sum_{a\in{\mathcal{A}}}\sum_{y\in\{0,1\}}\frac{\mathbb{P}[f=0\mid A=a]}{q_{y\mid a}\varepsilon^{2}})$.
@@
Note that $N_{y\mid a}$ samples are collected by the algorithm under the conditional distribution $\mathbb{P}[Y\mid A=a]$. Therefore, the expected number of samples actually requested by the algorithm is $N_{y\mid a}\cdot\mathbb{P}[f=0\mid A=a]$. Similarly, the audit cost of $N_{y\mid a}$ samples is $N_{y\mid a}(c_{feat}\mathbb{P}[f=0\mid A=a]+c_{lab}\mathbb{P}[f=0,Y=0\mid A=a])$. Therefore, the expected number of samples requested by the algorithm is at most
$\sum_{y\in\{0,1\}}\sum_{a\in{\mathcal{A}}}N_{y\mid a}\cdot\mathbb{P}[f=0\mid A=a]\leq\sum_{a\in\mathcal{A}}\sum_{y\in\{0,1\}}\frac{624\cdot\mathbb{P}[f=0\mid A=a]\log(8\lvert\mathcal{A}\rvert/\delta)}{\varepsilon^{2}q_{y\mid a}}~.$
@@
cost $=\sum_{y\in\{0,1\}}\sum_{a\in{\mathcal{A}}}N_{y\mid a}(c_{feat}\mathbb{P}[f=0\mid A=a]+c_{lab}\mathbb{P}[f=0,Y=0\mid A=a])$
$\leq\sum_{a\in\mathcal{A}}\sum_{y\in\{0,1\}}\frac{624(c_{feat}\mathbb{P}[f=0\mid A=a]+c_{lab}\mathbb{P}[f=0,Y=0\mid A=a])\log(8\lvert\mathcal{A}\rvert/\delta)}{\varepsilon^{2}q_{y\mid a}}~.$
@@
Suppose the true hypothesis is $\mathsf{UNFAIR}$. Let $y^{*}\in\{0,1\}$ and $a_{1},a_{2}\in\mathcal{A},a_{1}\neq a_{2}$ be such that $\lvert\mathbb{P}[f=1\mid Y=y^{*},A=a_{1}]-\mathbb{P}[f=1\mid Y=y^{*},A=a_{2}]\rvert=\left\lvert\frac{p_{y^{*}\mid a_{1}}}{q_{y^{*}\mid a_{1}}}-\frac{p_{y^{*}\mid a_{2}}}{q_{y^{*}\mid a_{2}}}\right\rvert>\gamma+\varepsilon$. Then, we have,
@@
$>\gamma+\varepsilon-\frac{\varepsilon}{2}$ ($\left\lvert\frac{p_{y^{*}\mid a_{1}}}{q_{y^{*}\mid a_{1}}}-\frac{p_{y^{*}\mid a_{2}}}{q_{y^{*}\mid a_{2}}}\right\rvert>\gamma+\varepsilon$)
$=\gamma+\frac{\varepsilon}{2}~.$
@@
$\lvert\mathbb{P}[f=1\mid Y=y,A=a]-\mathbb{P}[f=1\mid Y=y,A=a^{\prime}]\rvert=\left\lvert\frac{p_{y\mid a}}{q_{y\mid a}}-\frac{p_{y\mid a^{\prime}}}{q_{y\mid a^{\prime}}}\right\rvert\leq\gamma~.$
@@
$\leq\gamma+\frac{\varepsilon}{2}~.$ ($\left\lvert\frac{p_{y\mid a}}{q_{y\mid a}}-\frac{p_{y\mid a^{\prime}}}{q_{y\mid a^{\prime}}}\right\rvert\leq\gamma$)
@@ A.3 Lower Bound for the Blackbox Model
Under the blackbox setting, for any $\varepsilon\in(0,\nicefrac{{1}}{{8}})$ and $\delta\in(0,1)$, there exists a family of instances characterized by $\gamma=0$, $q_{y\mid a}$, and $\mathbb{P}[f=0\mid A=a]$ for $y\in\{0,1\}$ and $a\in{\mathcal{A}}$, even with $\lvert\mathcal{A}\rvert=2$, such that any algorithm with finite expected running time that succeeds in the $(\varepsilon,\delta)$-fairness test must request at least $\widetilde{\Omega}\left(\sum_{a\in\mathcal{A}}\sum_{y\in\{0,1\}}\frac{\beta_{a}}{q_{y\mid a}\varepsilon^{2}}\right)$ true labels in expectation.
@@
Let $p,q\in(0,1)$ be constants to be decided later. We let $\mathbb{P}[Y=1\mid A=a]=q$. The classifier $f$ is given as follows. Under the $\mathsf{FAIR}$ hypothesis, given $Y=1$ and $A=a$, $f\sim\mathrm{Ber}(\frac{1}{2})$ and given $Y=0$ and $A=a$, $f\sim\mathrm{Ber}(\frac{p}{2(1-q)})$ for both $a\in\{0,1\}$. On the other hand, in the $\mathsf{UNFAIR}$ hypothesis, the distribution of $f$ is as specified in Table 1.
@@
From Table 2, it can be seen that given $A=a$, $\mathbb{P}_{\mathsf{FAIR}}[f=1\mid A=a]=\mathbb{P}_{\mathsf{UNFAIR}}[f=1\mid A=a]$. Therefore, $\mathbb{P}_{\mathsf{FAIR}}[f=0\mid A=a]=\mathbb{P}_{\mathsf{UNFAIR}}[f=0\mid A=a]=1-\frac{p+q}{2}$ for all $a\in cA$. Moreover, the $\mathbb{P}_{\mathsf{FAIR}}[Y=1\mid f=1,A=a]=\mathbb{P}_{\mathsf{UNFAIR}}[Y=1\mid f=1,A=a]$. Hence, $D_{KL}(\mathbb{P}_{\mathsf{FAIR}}[Y\mid f=1,A=a]\parallel\mathbb{P}_{\mathsf{UNFAIR}}[Y\mid f=1,A=a])=0$. Finally, $\mathbb{P}_{\mathsf{FAIR}}[Y=1\mid f=0,A=1]=\frac{q}{2(1-\frac{p+q}{2})}$ and $\mathbb{P}_{\mathsf{UNFAIR}}[Y=1\mid f=0,A=1]=\frac{q(1+8\varepsilon)}{2(1-\frac{p+q}{2})}$.
We set $\varepsilon\in(0,\frac{1}{8})$, $p\in(0,\frac{1}{3})$ and $q\in(0,\frac{1}{3})$ so that $q(1+4\varepsilon)\leq 1$. It is easy to verify that the equalized odds difference under the $\mathsf{FAIR}$ hypothesis is 0 (intuitively, the probabilities of the classifier is the same for both the groups $a\in\{0,1\}$). The equalized odds disparity gap of the $\mathsf{UNFAIR}$ hypothesis is more than $\varepsilon$ as can be seen below.
@@
$\geq\frac{1}{2}\exp\left(-64\cdot\mathbb{E}_{\mathsf{FAIR}}[L_{\tau}]\cdot\varepsilon^{2}\cdot\frac{q}{2(1-\frac{p+q}{2})}\right)$ (Lemma 12, $\frac{q}{2-p-q}\leq\frac{1}{3}$ and $\frac{8q\varepsilon}{2-p-q}\leq\frac{1}{2}$ for $p,q\in(0,\frac{1}{3})$ and $\varepsilon\in(0,\frac{1}{4})$)
Therefore, $\mathbb{E}_{\mathsf{FAIR}}[L_{\tau}]>\frac{(1-\frac{p+q}{2})\log(1/4\delta)}{32q\varepsilon^{2}}$. Moreover, since Lemma 11 is symmetric, we also have $\mathbb{E}_{\mathsf{UNFAIR}}[L_{\tau}]>\frac{(1-\frac{p+q}{2})\log(1/4\delta)}{64q\varepsilon^{2}}$ using Lemma 12. Hence, $\min\{\mathbb{E}_{\mathsf{FAIR}}[L_{\tau}],\mathbb{E}_{\mathsf{UNFAIR}}[L_{\tau}]\}>\frac{(1-\frac{p+q}{2})\log(1/4\delta)}{64q\varepsilon^{2}}$. Therefore, under both the hypothesis, the algorithm must request $\Omega\left(\frac{\mathbb{P}[f=0\mid A=1]\log(1/\delta)}{\mathbb{P}[Y=1\mid A=1]\varepsilon^{2}}\right)$ true labels in expectation to be able to output the correct hypothesis. Further, each requested true label is associated with a cost of $c_{feat}+c_{lab}\mathbb{P}[Y=0\mid f=0]$. Thus, the total cost is at least $\Omega\left(\frac{(c_{feat}+c_{lab}\mathbb{P}[Y=0\mid f=0])\cdot\mathbb{P}[f=0\mid A=1]\log(1/\delta)}{\mathbb{P}[Y=1\mid A=1]\varepsilon^{2}}\right)=\Omega\left(\frac{(c_{feat}\mathbb{P}[f=0\mid A=1]+c_{lab}\mathbb{P}[f=0,Y=0\mid A=a])\log(1/\delta)}{\varepsilon^{2}}\right)$.